If roberts age is 3/4 of his brothers age and robert is 45 years old. How old is his brother?
Answer by oscargut(2103)

Let x = brotherīs age
(3/4)x = 45
x = 45(4/3) = 60
Answer: 60
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com